OpenStudy (anonymous):

in triangle ABC, E is the midpoint of AC and D is the midpoint of CB. if DF is parallel to BE, find the length of side AB.

OpenStudy (anonymous):

Is there an image?

jhonyy9 (jhonyy9):

us proportionalitie of sides

jhonyy9 (jhonyy9):

first try drawing it

OpenStudy (anonymous):

 here is the figure

jhonyy9 (jhonyy9):

so if D is the midpoint of CB what will be the lenght of CB ? this is easy

OpenStudy (anonymous):

10? i'm not sure

jhonyy9 (jhonyy9):

yes you are right

jhonyy9 (jhonyy9):

so then you know CB=10

OpenStudy (anonymous):

then?

jhonyy9 (jhonyy9):

now you can writing the proportionalitie of sides CD CB ---- = ---- FD BF and from this you can calculi easy the BF ok?

OpenStudy (anonymous):

BF= 6

jhonyy9 (jhonyy9):

5 10 --- = --- 3 BF cross multiplie 5*BF = 3*10 yes you are right

OpenStudy (anonymous):

what's next after getting BF?

jhonyy9 (jhonyy9):

the same you can calculi CE

jhonyy9 (jhonyy9):

CF FD --- = ---- CE EB

OpenStudy (anonymous):

CE- 8 am i correct?

jhonyy9 (jhonyy9):

then CE= (EB*CF)/FD

jhonyy9 (jhonyy9):

sorry BE=6 not BF BF not exist

jhonyy9 (jhonyy9):

yes CE=8

jhonyy9 (jhonyy9):

so E is the midpoint of AC what mean AC= ?

jhonyy9 (jhonyy9):

then the lenght of AC = ?

OpenStudy (anonymous):

AC= 16

OpenStudy (anonymous):

then AB = 10

OpenStudy (anonymous):

is it correct?

jhonyy9 (jhonyy9):

how you have got AB=10 ?

OpenStudy (anonymous):

\[\left(\begin{matrix}CE \\ AC\end{matrix}\right)= \left(\begin{matrix}FD \\ BE\end{matrix}\right)\]

OpenStudy (anonymous):

then by pythagorean theorem..

jhonyy9 (jhonyy9):

ok i understand it now you are right

OpenStudy (anonymous):

Yaay.. thanks a lot :)

jhonyy9 (jhonyy9):

my pleasure good luck